From GlueXWiki
Jump to: navigation, search

September 11, 2014 FDC+CDC meeting


  1. DAQ (Beni)
  2. CDC update (Mike)
  3. FDC update (Vlad, Lubomir)
  4. Electronics (Fernando)
  5. Engineering
  6. Commissioning Wiki Page (all)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Naomi (CMU), Fernando, Beni, Eugene, Mike, Bill, Nick, Simon, Fernando, Vlad, and Lubomir (JLab).


- Beni: new fADC125 firmware installed. First tests of the CDC crates: all OK except one module in crate 3. Beni will modify his code to accept new data types. Hopefully he will start run tomorrow using mode 8: raw data, time, and integral. For that one needs to define proper thresholds (several sigmas of the noise above pedestal), and daq values.

- Record long run, #304, has been take with all the CDC and 12 (out of 14) FDC crates. The run didn't crash after >5 days, we had to stop it for the new firmware to be installed.

- After the power outage yesterday (at ~1:30pm) Beni had to go down in the hall to turn on the VME crates, including the third trigger crate. One should be able to do this from the web, it's not clear why it didn't work.

CDC update

- Mike: few pre-amps have been changed. Four fADC125 modules were replaced due to missing bits in some channels. On the next meeting Mike will present results about the straw efficiency.

- Due to the power outage the mass flow controller stopped and the pressure in the CDC went down to atmospheric in about 10-20 minutes. Beni: need to check for gas leaks with sniffer.

FDC update

- In run #304 two crates were off: one due to a new firmware installed, the other had issues with one module. In addition we had 6 modules not reporting any signals, and many channels with pedestals below zero. We have to understand what's wrong with these 6 modules, and replace them if we have to.

- Vlad: looked into the run 304 results, for the first time we have enough statistics (at least 10 events in each channel) to decide if the channel is working. Results: 1 wire channel (3M connector problem) and 16 strip channel not working. The low number of bad strips (to be compared to 60 bad channels before) is due the low threshold used in the analyses. However, 4 of the 16 bad channels are new.

- Due to the power outage the chiller, the LV and the gas in the Hall went off. We turned on the chiller successfully, even it took ~20 min till it reached the 20deg target temperature. As for the gas it was restored, but we need to connect the two mass flow controller boxes to the UPS (we will need extension cords and Tom's permission). The two gas boxes in the gas room were connected to a UPS and we had no problems there. There was a problem to turn the LV back on. Even the chiller was on and the hardware interlock not engaged, we couldn't do it yesterday. Today, according to Hovanes, it turned out that there's another IOC (independent of the FDC IOC) that checks the temperature and the flow in the chiller and interlocks (using epics software) the LV crates. In the mean time Dave Butler made the status of the hardware interlock available for epics. Thus, Hovanes is working now on a control window that will have all this information so that one can monitor the interlocks and clear them to be able to start the LV.


- Fernando: there are cables in the Hall that are not marked, run properly, or even have fire hazard (mostly in the FCAL area). Related to the FDC we have a cable from the chiller (used for the epics) and an ethernet cable that has to be marked and run properly.

- Fernando and Nick installed ferrite core filters on all the HV cables (CDC and FDC) at the module side, but the noise didn't change. The plan is to do nothing for the coming run. Later will consider rearranging the HV cables in the CDC.


- Linked above are the very last Luke's results (we didn't have time to discuss them on the last meeting). Luke re-measured the first position. The second plots shows the difference between the new and the old pitch values: the problem is that there's a constant slope. This is well seen from the third plot, showing the difference between the new and the old measurements point by point (i.e. the second plot is the derivative of the third one). The difference goes up to 0.3mm in the middle. The new measurement was done when the AC was off in the ESB, so the discrepancy might be related to the humidity/temperature change.


- Beni: not sure if using the Bertan modules on some of the CDC channels is possible, since there's no way to control them remotely, after the beam is aligned it will require going in the hall to turn them on or to adjust the HV or trip levels.

- The CDC instructions for the initial detector checkout need to be written as line-by-line instructions for the shifters (according to Alex).